OpenStudy (anonymous):

Find all values of theta where 0 degrees < 0 < 360 degrees. a. csc(theta) = - √2 b. sin(theta) = -√3/2 c. sec(theta) = -1

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

Nothing.. Completely lost and I need to finish this asap.

jimthompson5910 (jim_thompson5910):

csc(theta) = 1/sin(theta)

jimthompson5910 (jim_thompson5910):

csc(theta) = - √2 1/sin(theta) = - √2 sin(theta) = - 1/√2 sin(theta) = -(√2 )/2 now use the unit circle to determine theta

jimthompson5910 (jim_thompson5910):

I'll let you take over. Tell me what you get

OpenStudy (anonymous):

Um.. 4pi/3?

jimthompson5910 (jim_thompson5910):

close but no

OpenStudy (anonymous):

But -1/2 , -√2, 2 isn't stated on the unit circle. I'm unsure how to find the measurement.

jimthompson5910 (jim_thompson5910):

try looking for \(\large -\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

That's in quadrants 3 & 4.

jimthompson5910 (jim_thompson5910):

yes because sine is negative here

jimthompson5910 (jim_thompson5910):

which angles correspond to a point with a y coordinate of \(\large -\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

225 , 315

jimthompson5910 (jim_thompson5910):

very good, you got them both

jimthompson5910 (jim_thompson5910):

so that is the solution to part A

OpenStudy (anonymous):

Oh okay that makes much more sense!

jimthompson5910 (jim_thompson5910):

you are doing the same for each part note: sec(theta) = 1/cos(theta)

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

Can you set the problems up again? Still confusing in that area.

jimthompson5910 (jim_thompson5910):

which one

OpenStudy (anonymous):

B

jimthompson5910 (jim_thompson5910):

in part B you need to look at the unit circle for points that have y coordinates of \(\large -\frac{\sqrt{3}}{2}\)

jimthompson5910 (jim_thompson5910):

then report the angles (in degrees) for those corresponding points

jimthompson5910 (jim_thompson5910):

remember, sine deals with the y coordinate on the unit circle

OpenStudy (anonymous):

240 , 300?

jimthompson5910 (jim_thompson5910):

both are perfect

OpenStudy (anonymous):

Awesome, what about C?

jimthompson5910 (jim_thompson5910):

sec(theta) = -1 1/cos(theta) = -1 cos(theta) = -1 now use the unit circle...BUT this time you're looking at the x coordinate (not y coordinate)

OpenStudy (anonymous):

180, ?

jimthompson5910 (jim_thompson5910):

and that's it, just 180

OpenStudy (anonymous):

Awesome, thank you so much! Do you mind helping me out with coterminals?

jimthompson5910 (jim_thompson5910):

sure I can do a few more

OpenStudy (anonymous):

Find two angles, theta, coterminal with -25pi/8 where -2pi<theta<2pi.

jimthompson5910 (jim_thompson5910):

use a calculator to evaluate -25pi/8 what do you get

OpenStudy (anonymous):

-9.817

jimthompson5910 (jim_thompson5910):

notice how -2pi = -6.28 roughly

jimthompson5910 (jim_thompson5910):

and 2pi = 6.28 roughly

jimthompson5910 (jim_thompson5910):

so clearly -25pi/8 is not in the interval -2pi<theta<2pi

OpenStudy (anonymous):

Okay

jimthompson5910 (jim_thompson5910):

to find any coterminal angles to a given angle (in radians) we just add/subtract 2pi

jimthompson5910 (jim_thompson5910):

in this case, we add 2pi

jimthompson5910 (jim_thompson5910):

-25pi/8 + 2pi = ??

OpenStudy (anonymous):

Um.. -23pi/8?

jimthompson5910 (jim_thompson5910):

no, 2pi = 4pi/2

OpenStudy (anonymous):

Oh one sec

jimthompson5910 (jim_thompson5910):

sorry I meant to say 2pi = 16pi/8

jimthompson5910 (jim_thompson5910):

-25pi/8 + 2pi -25pi/8 + 16pi/8 = ??

OpenStudy (anonymous):

-9pi/8?

jimthompson5910 (jim_thompson5910):

now evaluate that with a calculator

OpenStudy (anonymous):

-3.53

jimthompson5910 (jim_thompson5910):

is that in the interval -6.28 < theta < 6.28 ??

OpenStudy (anonymous):

Yes it is :)

OpenStudy (anonymous):

What about the other one?

jimthompson5910 (jim_thompson5910):

so that's one coterminal angle that's in the right interval we want add 2pi to -9pi/8 to get the next coterminal angle

OpenStudy (anonymous):

7pi/8?

jimthompson5910 (jim_thompson5910):

very good, you got the hang of this

OpenStudy (anonymous):

2.74

jimthompson5910 (jim_thompson5910):

so that's definitely in range too

jimthompson5910 (jim_thompson5910):

the two coterminal angles are -9pi/8, 7pi/8

OpenStudy (anonymous):

Great thanks! I have 2 more different problems to solve rather quickly because I can figure the rest off the bases of these. Do you mind? You're a great help.

jimthompson5910 (jim_thompson5910):

sure go for it

OpenStudy (anonymous):

Find the measure of the reference angle: a. theta = 18pi/7 Find the radian measure of the angle: theta = -270 degree measure of : theta=8pi/12 Sorry 3 actually.

jimthompson5910 (jim_thompson5910):

evaluate 18pi/7 with a calculator

OpenStudy (anonymous):

8.07

jimthompson5910 (jim_thompson5910):

so that's definitely out of range of 0 < theta < 2pi

jimthompson5910 (jim_thompson5910):

so we need to find the coterminal angle subtract 2pi from 18pi/7 to get 18pi/7 - 2pi = ???

OpenStudy (anonymous):

Hmm the denominators are different..

jimthompson5910 (jim_thompson5910):

2pi = 14pi/7

jimthompson5910 (jim_thompson5910):

multiply top and bottom by 7

OpenStudy (anonymous):

Oh okay. So 4pi/7.

jimthompson5910 (jim_thompson5910):

4pi/7 = 1.795 so 4pi/7 is definitely in the interval 0 < theta < 2pi

OpenStudy (anonymous):

Oh wait. This is to find the reference angle.

jimthompson5910 (jim_thompson5910):

the question is: which quadrant is 4pi/7 in? well we could do a bunch of computations of pi/2, pi, 3pi/2 and compare, or we can convert 4pi/7 to degrees

jimthompson5910 (jim_thompson5910):

I would convert it over to degrees so you multiply this by 180/pi to get (4pi/7)*(180/pi) = (4pi*180)/(7*pi) = 720/7 = 102.857 now what quadrant is the angle 102.857 degrees in?

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so we use this rule Reference Angle = pi - (Given Angle)

jimthompson5910 (jim_thompson5910):

Reference Angle = pi - (Given Angle) Reference Angle = pi - 4pi/7 Reference Angle = 7pi/7 - 4pi/7 Reference Angle = 3pi/7

OpenStudy (anonymous):

Makes sense! Awesome!

jimthompson5910 (jim_thompson5910):

ok that's great

OpenStudy (anonymous):

Now for question 2

jimthompson5910 (jim_thompson5910):

Find the radian measure of the angle: theta = -270

jimthompson5910 (jim_thompson5910):

multiply this by pi/180 and reduce

OpenStudy (anonymous):

-540?

jimthompson5910 (jim_thompson5910):

no -270*(pi/180) = ??

OpenStudy (anonymous):

-270pi/-1.5 ? Er.

jimthompson5910 (jim_thompson5910):

-270*(pi/180) = -270pi/180 which reduces to what?

OpenStudy (anonymous):

-4.7

jimthompson5910 (jim_thompson5910):

leave it as a fraction though

jimthompson5910 (jim_thompson5910):

it might help to just focus on reducing -270/180

OpenStudy (anonymous):

-9pi/4?

jimthompson5910 (jim_thompson5910):

close, it should be -3pi/2

jimthompson5910 (jim_thompson5910):

now add 2pi to this

OpenStudy (anonymous):

pi/2?

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

:) Question 3 and I'll leave ya alone, :)

jimthompson5910 (jim_thompson5910):

degree measure of : theta=8pi/12

jimthompson5910 (jim_thompson5910):

multiply this by 180/pi and reduce